The Thomas Supply Company Inc. is a distributor of gas-powered generators. As with any business, the length of time customers take to pay their invoices is important. Listed below, arranged from smallest to largest, is the time, in days, for a sample of The Thomas Supply Company Inc. invoices.

   13

13

13

20

26

29

32

34

34

34

35

35

36

37

38

   41

41

41

45

45

47

47

49

53

54

55

56

62

67

82

(Round your answers to 2 decimal places.)
a. Determine the first and third quartiles.
Q1 =
Q3 =
b. Determine the second decile and the eighth decile.
D2 =
D8 =
c. Determine the 67th percentile.
Answer #1

Solution:

Given that,

Arranging Observations in the ascending order, We get :
13,13,13,20,26,29,32,34,34,34,35,35,36,37,38,41,41,41,45,45,47,47,49,53,54,55,56,62,67,82

Here, n = 30

a ) Q1 = ( n + 1 / 4 )th Value of observation.

= ( 31/ 4 )th Value of observation.

   = 7.75th Value of observation.

= 7th observation. + 0.75 (8th - 7th )

= 32 +0.75 ( 34 - 32 )

= 32 + 0.75 (2 )

= 32 + 1.5

= 33.5

Q1 = 33.5

Q3 = ( 3 (n + 1) / 4 )th Value of observation.

= ( 31/ 4 )th Value of observation.

   = 23.25th Value of observation.

= 23th observation. + 0.25 (24th - 23th )

= 49 +0.25 ( 53 - 49 )

= 49 + 0.25 (4 )

= 49 + 1

= 50

Q3 = 50

b) D2 = ( 2 (n + 1) / 10 )th Value of observation.

= ( 2*31 / 10 )th Value of observation.

   = 6.20th Value of observation.

= 6th observation. + 0.20 (7th - 6th )

= 29 +0.20 ( 32 - 29 )

= 29 + 0.20 (3 )

= 29 + 0.60

= 29.6

D2 =29.6

D8 = ( 8 (n + 1) / 10 )th Value of observation.

= ( 8*31 / 10 )th Value of observation.

   = 24.8th Value of observation.

= 24th observation. + 0.8 (25th - 24th )

= 53 +0.8 ( 54 - 53 )

= 53 + 0.8 (1 )

= 53 + 0.8

= 53.8

D8 = 53.8

C) P67 = ( 67 (n + 1) / 100 )th Value of observation.

= ( 67*31 / 100 )th Value of observation.

   = 20.77th Value of observation.

= 20th observation. + 0.77 (21th - 20th )

= 45 +0.77 ( 47 - 45 )

= 45 + 0.77 (2)

= 45 + 1.54

= 46.54

P67 = 46.54